# Recommender systems and reinforcement learning for building control and occupant interaction: A text-mining driven review of scientific literature

### v1

@@ 2.4 Step 4: NLP text mining using Word2Vec
Word2Vec is a Python-based word embedding tool that is used to identify semantic similarities between words within a text [64, 65]. This method maps each word to a multidimensional vector, facilitating the measurement of similarity through the cosine similarity of these vectors. Words with vectors that are closer in angle are considered to be more semantically similar.
The objective of training with Word2Vec is to identify an optimal vector representation for each word in the vector space that reflects its semantic meaning. Specifically, Word2Vec utilizes context within sentences, defined as a window that encompasses n words surrounding the target word, to predict the target word or the center word. During training, each word in a sentence is sequentially designated as the center word. The vector representations for both the center word and its surrounding context words are updated based on the discrepancy between the predicted center word vector and its actual vector. Training continues until these discrepancies no longer significantly decrease or model convergence. Convergence implies that the vectors calculated based on surrounding words are sufficiently close to the actual vector of the center word or that further training does not significantly enhance model performance. In this study, a context window of 20 words was used, and each word was assigned a 300-dimensional vector, settings that were determined by hyperparameter tuning.
@@ 2.5 Step 5: Extracting the relationship between categories
In the Word2Vec model, the similarity between two words is determined by the frequency of their co-occurrence within similar contexts. For example, if the words thermal_comfort and setpoint appear frequently along with comparable terms, they are represented by vectors that are relatively close in the vector space. The trained word2vec model is used to extract similar words and the similarity between each word in a different category.
An initial step involves the comprehensive extraction of relevant words from each category. Predefined words within each category are used as input to the NLP model to extract words that are semantically similar. This procedure aims to identify and retrieve all relevant words associated with each category from the existing literature. For each word, the model extracts the top 100 similar words. Subsequently, any terms deemed irrelevant are removed using the GPT-4 model. This step ensures that the final dataset is refined and relevant to the specific categories under study. After this preliminary filtering, any remaining words are subjected to manual review and further refinement by human analysts.
@@
Just-In-Time Adaptive Interventions (JITAI) A system designed to provide intervention on an as-needed basis, adapting in real time to an individual’s changing circumstances and conditions [85].
@@
3 Results
In this analysis, heat maps are utilized to demonstrate the relevance between pairs of words. Additionally, hierarchical clustering is used to form sub-clusters within each category, which serve to present the compositional structure of each category and enhance the analysis of interrelationships among these clusters. Specifically, Hierarchical Agglomerative Clustering (HAC) utilizing Ward’s method was applied to the word embedding vectors within each category. This approach clusters similar words based on their Euclidean distances in the vector space, facilitating the representation of these clusters in a tree structure known as a dendrogram. Figures 4 and 5 illustrate the data mining results and subclusters within the categories of input data, objectives, algorithms, and platforms.
@@ 3 Results
3.1 Input data and objectives of recommender systems for building control and occupant interaction
The initial category extracted the strength of relationships between various input data/objectives and different types of recommender systems. The horizontal axis (Recommender Systems) categorizes the types of recommender systems commonly used in building control and occupant interaction, as identified in the literature. The vertical axis (Input Data/Output Objectives) is organized by categories of terms. The left diagram presents the average relevance of each category of output terms to the recommender systems, while the right diagram provides a detailed classification of outputs.
Figure 6 highlights the significant relevance of real-time and cloud-based data (DT) in this domain, which underscores the sector’s high demand for real-time data, with the majority of recommender system algorithms computed on cloud platforms [10]. The most pertinent data categories for these recommender systems are user context and sensor data (CS). These data types are crucial, as they directly enhance the personalization of recommendations, leveraging real-time environmental and user-specific information to optimize system output. JITAI particularly benefits from high correlations with physiological data (PD) and ecological momentary assessment to provide interventions related to indoor health and behavioral modifications. However, data related to indoor air quality and comfort (IAQ & ITC), such as indoor pollutant concentrations, carbon dioxide levels, ventilation rates, indoor temperatures, noise levels, predicted percentage of dissatisfied occupants, and clothing insulation factor, exhibit lower correlations with recommender systems in the built environment. These data types, despite being crucial for the built environment, are not sufficiently utilized in current recommender systems.
Figure 7 provides significant insight into how different types of recommender system align with objectives pertinent to the control of the building and the interaction of the occupants. Regarding the types of recommender systems, it highlights that JITAI is the most widely applicable form of recommendation system within this domain, particularly effective in delivering healthcare and behavioral change interventions (HL). However, reinforcement learning-based recommender systems are not as widely used (showing the lowest relevance to the objectives), but are more closely associated with interventions in indoor environmental comfort and energy performance (IEE). This limited application may be due to the complexity involved in these areas. Regarding the objectives of recommender systems, specific content recommendations (SC), which include suggestions on music, routes, and locations, demonstrate strong associations across all types of recommender systems. However, the analysis reveals a significant gap in the application of recommender systems to objectives directly related to indoor environmental comfort and energy efficiency (IEE). Although these are primary objectives of building control, recommender systems are currently underutilized in these areas.
@@ 3.2 Input data for different recommendation objectives
The subsequent comparison provides an overview of the relationship between input data and output in recommender systems, focusing solely on understanding the types of input data required for different outputs. In this analysis, both axes are organized based on the average strength of relationships, with the horizontal axis representing the objectives of the recommender systems (arranged from left to right) and the vertical axis detailing the clusters of input data (arranged from top to bottom).
@@ 3.3 Algorithms used in recommender systems
The next comparison uses words related to the algorithm and the types of recommender systems, as shown in Figure 9. This heat map analysis reveals distinct patterns in the applicability of different algorithms across various types of recommender systems for building control and occupant interaction. Predominantly used algorithms include collaborative filtering, content-based, knowledge-driven, fusion-based, and clustering algorithms. NLP exhibits significant effectiveness in content-based and knowledge-driven systems due to its capacity to analyze and interpret complex user data and contextual information. In contrast, regression and optimization algorithms demonstrate a weaker association with most types of recommender systems examined. This may suggest that these more traditional algorithms are less favored for complex or dynamic recommendation tasks that require adaptation to new information or user behaviors. Moreover, algorithms such as reinforcement learning display variable degrees of relationship strength across different types of recommender system, suggesting their applicative versatility and potential underexplored areas for their application.
@@ 4 Discussion
Given the prominence of recommender systems in health management and their growing relevance in the mobile health sector, substantial research has already been published in medical and mobile health journals. Therefore, the subsequent sections will focus on discussing the application of recommender systems in building control and occupant interaction to elucidate their potential and identify gaps in this domain. Specifically, this section will address the following three questions:
@@
Figure 11 presents the relationships between different types of algorithms and the various applications of recommender systems, with a specific emphasis on the application of each algorithm in the field of building control and occupant interaction. Both axes are organized based on the strength of relationships, with the horizontal axis representing the algorithms (arranged from left to right) and the vertical axis detailing the clusters of objectives (arranged from top to bottom). The following subsection will, based on the insights derived from this figure, elucidate key considerations that the research community can address in future studies.
@@
Figure 11: The figure on the left shows the relation between algorithms and recommendation objectives sorted based on usability. The figure on the right shows the relation between algorithms and built environment applications sorted based on the group correlation and element correlation.
@@
4.1 Algorithm selection for applications in building control and occupant interaction
@@
Based on the left side of Figure 11, it can be seen that the algorithms most commonly used for recommender systems include traditional methods such as content-based, collaborative filtering, and knowledge-based systems. However, for applications in building control and occupant interaction, where the variety of data types is extensive, the most commonly used recommendation algorithms are predominantly based on learning algorithms and traditional ML techniques. The five most commonly used recommendation algorithms, illustrated on the right side of Figure 11 (highlighted with black borders), are reinforcement learning, knowledge-based, clustering, fusion-based, and deep learning.
@@
Reinforcement learning-based algorithm operates on the principle of learning optimal actions through interactions with the environment to maximize cumulative rewards [84]. These algorithms continually adjust their strategies based on user interaction feedback, refining their recommendations over time [86]. This adaptive capability makes RL-based algorithms particularly suitable for applications in the built environment, where user preferences, environmental conditions, and energy consumption patterns frequently change [6]. In such applications, value-based methods like Deep Q-Network (DQN) and Double Deep Q-Network (DDQN) algorithms are extensively utilized for energy performance optimization [87] and recommender system design [88]. Owing to their online learning capabilities, which include real-time strategy updates and experience replay mechanisms, these algorithms ensure that the models remain current and efficient [89]. Furthermore, RL-based algorithms are well-suited for addressing multi-objective optimization problems in complex environments. For instance, by integrating XGBoost with DQN, researchers have optimized indoor thermal comfort and energy consumption by predicting the impact of occupants’ behaviors on indoor environments [5]. Despite these advantages, RL algorithms face challenges such as the cold start problem, which is when the system lacks sufficient initial data to make accurate recommendations [38], time-intensive training processes [4], and poor interpretability [89]. Addressing the cold start issue requires substantial data for model training or the incorporation of prior knowledge during the design phase [90, 91]. Moreover, techniques like imitation learning and transfer learning are employed to reduce training times [92, 93], while [94] employs explainable AI to address the challenge of poor interpretability.
@@
Knowledge-based algorithms leverage domain-specific knowledge to generate recommendations, which is crucial when explicit user preferences are unavailable, and decisions must be guided by objective standards or expert knowledge [81]. In the context of the built environment, these algorithms can utilize detailed knowledge about building operations, environmental parameters, and user preferences to recommend optimal strategies [10, 25, 95]. For example, a knowledge-based algorithm can be programmed to maintain acceptable indoor pollutant concentration levels. When sensors detect pollutant levels exceeding predefined thresholds, the system recommends increased ventilation to maintain air quality [20]. Moreover, knowledge-based approaches are often used to enhance other algorithms. For instance, the HeartSteps algorithm, developed to promote physical activity, integrates reinforcement learning with prior knowledge derived from historical data to overcome the cold start problem typical of traditional recommender systems [12].
@@
Clustering algorithms enhance targeted and efficient recommendations by grouping similar items or users based on their attributes or behavioral characteristics using clustering techniques [96]. Unlike collaborative filtering, which relies on user ratings or behavioral similarity [97], clustering algorithms focus on identifying the inherent grouping structure within the data [98, 99], thereby significantly mitigating the cold start problem commonly encountered in collaborative filtering due to data sparsity [100]. For instance, clustering methods like K-means [101] and DBSCAN [102] can perform extensive computations during the preprocessing stage, thereby reducing the need for real-time calculations and improving the operational efficiency of recommender systems. Despite their advantages in handling large-scale datasets, clustering algorithms exhibit limitations in dynamic environments, such as building control or indoor environment optimization [103]. Clustering algorithms typically require major computations, such as determining cluster centers and assigning data points to clusters, to be completed during the preprocessing phase [96, 104]. This means that whenever there are changes in the dataset, such as the addition of building operational data, the clustering model may need to be rerun to reflect these changes. This process is not automatic, contrasting with the real-time responsiveness of collaborative filtering [97].
@@
Deep learning-based algorithms utilize complex neural network architectures to model user preferences and item characteristics within a recommendation framework [113]. These systems employ multi-layer nonlinear processing units to learn deep representations of features from complex user behaviors, item properties, and textual descriptions, enabling them to discern intricate patterns in large-scale data that traditional algorithms might overlook [114, 115]. For instance, in [116], deep learning models were used to address significant cold start issues in collaborative filtering systems by embedding information about users and their preferences directly into the model. Moreover, deep learning techniques are frequently applied to process unstructured data, such as text, images, and audio [10, 117]. Such as Convolutional Neural Networks (CNNs) [118, 119] are used for image recommendations, while Recurrent Neural Networks (RNNs) [120] handle sequential interaction data. Emerging technologies like attention mechanisms and Transformer models [121] have been widely adopted in deep learning-based recommender systems [114]. These models focus more on the most pertinent parts of the input data, dynamically adjusting the importance weights of different segments to optimize information processing and feature extraction [122]. For instance, in [123], the incorporation of time-aware self-attention mechanisms enhances the precision of capturing user preferences, significantly improving the performance of recommendation models.
4.2 Well-established applications of recommender systems in building control and occupant interaction

### v2

@@ 2.4 Step 4: NLP text mining using Word2Vec
Word2Vec is a Python-based word embedding tool that is used to identify semantic similarities between words within a text [79]. During the Word2Vec model training, words are projected into a vector space based on their semantic context within surrounding text. Words with similar meanings tend to be positioned closer to each other in this vector space. The relationships between these words are then quantified using the cosine similarity between their respective vectors (Equation 1).
$\text{sim}(A,B)=\cos(\theta)=\frac{A\cdot B}{\|A\|\|B\|}$ (1)
The objective of training with Word2Vec is to identify an optimal vector representation for each word in the vector space that reflects its semantic meaning. Specifically, Word2Vec utilizes context within sentences, defined as a window that encompasses n words surrounding the target word, to predict the target word or the center word. During training, each word in a sentence is sequentially designated as the center word. The vector representations for both the center word and its surrounding context words are updated based on the discrepancy between the predicted center word vector and its actual vector. Training continues until these discrepancies no longer significantly decrease or model convergence. Convergence implies that the vectors calculated based on surrounding words are sufficiently close to the actual vector of the center word or that further training does not significantly enhance model performance.
In this study, the preprocessed full texts of approximately 27,595 articles were used as training data for the Word2Vec model. Before training the model, hyperparameter configuration was conducted. Specifically, the min_count parameter was set to 2 to remove words with a frequency of less than 2, thereby reducing noise. To optimize the remaining hyperparameters, including vector_size and window, a synonym list containing 50 pairs of domain-specific terminology was created. This synonym list was developed based on a prior literature review. Hyperparameter tuning was conducted by evaluating the similarity between each pair of terms in the list under different hyperparameter combinations. The objective function was the average similarity of terms in the list(Equation 2), a higher average similarity indicated better model performance.
Objective $\text{Objective function}=\frac{1}{50}\sum_{i=1}^{50}sim(A_{i},B_{i})$ (2)
After testing 25 hyperparameter combinations, the optimal model was identified with a context window of 20 words and a vector size of 300, achieving a maximum overall similarity of 0.714. Furthermore, the study conducted analogy tests to evaluate whether the model successfully learned the semantic relationships between words. The principle behind these tests involves performing vector operations on words with specific semantic relationships to derive other words with analogous relationships. For example, in the model, the operation hvac -- thermostat $+$ lighting results in the vector corresponding to dimmer. This indicates that the model effectively captured the semantic relationship between the input words, as the relationship between hvac and thermostat parallels that between lighting and dimmer.
@@ 2.5 Step 5: Extracting the relationship between categories
After the training process, the trained Word2Vec model can be used to identify words related to the input word based on their proximity in the vector space, and compute the similarity between two input words by calculating the cosine similarity between their respective word vectors. Before extracting the correlation between two words, predefined words within each category are used as input to the NLP model to extract words that are semantically similar. This procedure aims to identify and retrieve all relevant words associated with each category from the existing literature. For each word, the model extracts the top 100 similar words. Subsequently, any terms deemed irrelevant are removed using the GPT-4 model. This step ensures that the final dataset is refined and relevant to the specific categories under study. After this preliminary filtering, any remaining words are subjected to manual review and further refinement by human analysts.
@@ 3 Results
Just-In-Time Adaptive Interventions (JITAI) A system designed to provide intervention on an as-needed basis, adapting in real-time to an individual’s changing circumstances and conditions [86].
@@
In this analysis, heat maps are utilized to demonstrate the relevance between pairs of words. The intensity of the heat map’s color represents the strength of the relationship: darker colors indicate stronger relationships, while lighter colors suggest weaker ones. These heat maps not only provide a quantitative representation of correlation but also serve as a tool to guide design decisions. For example, in heat map showing the relationship between input data and design objectives, the strength of the connections can help identify the necessary input data to achieve specific objectives, as well as irrelevant data types. Similarly, heat map depicting the relationship between recommendation system types and objectives can help determine the most suitable system type for a given design objective. Additionally, hierarchical clustering is used to form sub-clusters within each category, which serve to present the compositional structure of each category and enhance the analysis of interrelationships among these clusters. Specifically, Hierarchical Agglomerative Clustering (HAC) utilizing Ward’s method was applied to the word embedding vectors within each category. This approach clusters similar words based on their Euclidean distances in the vector space, facilitating the representation of these clusters in a tree structure known as a dendrogram. Figures 4 and 5 illustrate the data mining results and subclusters within the categories of input data, objectives, algorithms, and platforms.
@@ 3.1 Input data for different recommendation objectives
The initial comparison provides an overview of the relationship between input data and output in recommender systems, focusing solely on understanding the types of input data required for different outputs. In this analysis, both axes are organized based on the average strength of relationships, with the horizontal axis representing the objectives of the recommender systems (arranged from left to right) and the vertical axis detailing the clusters of input data (arranged from top to bottom).
@@
3.2 Input data and objectives of recommender systems for human-building interaction and occupant support
The subsequent category extracted the strength of relationships between various input data/objectives and different types of recommender systems. The horizontal axis (Recommender Systems) categorizes the types of recommender systems commonly used in human-building interaction and occupant support, as identified in the literature. The vertical axis (Input Data/Output Objectives) is organized by categories of terms. The left diagram presents the average relevance of each category of output terms to the recommender systems, while the right diagram provides a detailed classification of outputs.
Figure 7 highlights the significant relevance of real-time and cloud-based data (DT) in this domain, which underscores the sector’s high demand for real-time data, with the majority of recommender system algorithms computed on cloud platforms [12]. The most pertinent data categories for these recommender systems are user context and sensor data (CS). These data types are crucial, as they directly enhance the personalization of recommendations, leveraging real-time environmental and user-specific information to optimize system output. JITAI particularly benefits from high correlations with physiological data (PD) and ecological momentary assessment to provide interventions related to indoor health and behavioral modifications. However, data related to indoor air quality and comfort (IAQ & ITC), such as indoor pollutant concentrations, carbon dioxide levels, ventilation rates, indoor temperatures, noise levels, predicted percentage of dissatisfied occupants, and clothing insulation factor, exhibit lower correlations with recommender systems. These data types, despite being crucial for the indoor environment quality, are not sufficiently utilized in current recommender systems.
Figure 8 provides significant insight into how different types of recommender system align with objectives pertinent to the control of the building and the interaction of the occupants. Regarding the types of recommender systems, it highlights that JITAI is the most widely applicable form of recommendation system within this domain, particularly effective in delivering healthcare and behavioral change interventions (HL). However, RL-based recommender systems are not as widely used (showing the lowest relevance to the objectives), but are more closely associated with interventions in indoor environmental comfort and energy performance (IEE). This limited application may be due to the complexity involved in these areas. Regarding the objectives of recommender systems, specific content recommendations (SC), which include suggestions on music, routes, and locations, demonstrate strong associations across all types of recommender systems. However, the analysis reveals a significant gap in the application of recommender systems to objectives directly related to indoor environmental comfort and energy efficiency (IEE). Although these are primary objectives of current building design, recommender systems are currently underutilized in these areas.
@@ 3.3 Algorithms used in recommender systems
The next comparison uses words related to the algorithm and the types of recommender systems, as shown in Figure 9. This heat map analysis reveals distinct patterns in the applicability of different algorithms across various types of recommender systems for human-building interaction and occupant support. Predominantly used algorithms include collaborative filtering, content-based, knowledge-driven, fusion-based, and clustering algorithms. NLP exhibits significant effectiveness in content-based and knowledge-driven systems due to its capacity to analyze and interpret complex user data and contextual information. In contrast, regression and optimization algorithms demonstrate a weaker association with most types of recommender systems examined. This may suggest that these more traditional algorithms are less favored for complex or dynamic recommendation tasks that require adaptation to new information or user behaviors. Moreover, algorithms such as RL display variable degrees of relationship strength across different types of recommender system, suggesting their applicative versatility and potential underexplored areas for their application.
@@ 4 Discussion
Given the prominence of recommender systems in health management and their growing relevance in the mobile health sector, substantial research has already been published in medical and mobile health journals. Therefore, the subsequent sections will focus on discussing the application of recommender systems in human-building interaction and context-aware support to elucidate their potential and identify gaps in this domain. Specifically, this section will address the following three questions:
@@
  4. 4.
    What are the benefits and limitations of text mining-based literature review?
@@
Figure 11 presents the relationships between different types of algorithms and the various applications of recommender systems, with a specific emphasis on the application of each algorithm in the field of human-building interaction and occupant support. Both axes are organized based on the strength of relationships, with the horizontal axis representing the algorithms (arranged from left to right) and the vertical axis detailing the clusters of objectives (arranged from top to bottom). The following subsection will, based on the insights derived from this figure, elucidate key considerations that the research community can address in future studies.
@@
Figure 11: The figure on the left shows the relation between algorithms and recommendation objectives sorted based on usability. The figure on the right shows the relation between algorithms and human-building interaction applications sorted based on the group correlation and element correlation.
@@
4.1 Algorithm selection for applications in human-building interaction and occupant support
@@
Based on the left side of Figure 11, it can be seen that the algorithms most commonly used for recommender systems include traditional methods such as content-based, collaborative filtering, and knowledge-based systems. However, for applications in human-building interaction and occupant support, where the variety of data types is extensive, the most commonly used recommendation algorithms are predominantly based on learning algorithms and traditional ML techniques. The five most commonly used recommendation algorithms, illustrated on the right side of Figure 11 (highlighted with black borders), are RL, knowledge-based, clustering, fusion-based, and deep learning.
@@
Reinforcement learning-based algorithm operates on the principle of learning optimal actions through interactions with the environment to maximize cumulative rewards [85]. These algorithms continually adjust their strategies based on user interaction feedback, refining their recommendations over time [87]. This adaptive capability makes RL-based algorithms particularly suitable for applications in the human-building interaction, where user preferences, environmental conditions, and energy consumption patterns frequently change [8]. In such applications, value-based methods like Deep Q-Network (DQN) and Double Deep Q-Network (DDQN) algorithms are extensively utilized for energy performance optimization [88] and recommender system design [89]. Owing to their online learning capabilities, which include real-time strategy updates and experience replay mechanisms, these algorithms ensure that the models remain current and efficient [90]. Furthermore, RL-based algorithms are well-suited for addressing multi-objective optimization problems in complex environments. For instance, by integrating XGBoost with DQN, researchers have optimized indoor thermal comfort and energy consumption by predicting the impact of occupants’ behaviors on indoor environments [91]. Despite these advantages, RL algorithms face challenges such as the cold start problem, which is when the system lacks sufficient initial data to make accurate recommendations [92], time-intensive training processes [93], and poor interpretability [90]. Addressing the cold start issue requires substantial data for model training or the incorporation of prior knowledge during the design phase [94, 95]. Moreover, techniques like imitation learning and transfer learning are employed to reduce training times [96, 97], while [98] employs explainable AI to address the challenge of poor interpretability.
@@
Knowledge-based algorithms leverage domain-specific knowledge to generate recommendations, which is crucial when explicit user preferences are unavailable, and decisions must be guided by objective standards or expert knowledge [82]. In the context of human-building interaction, these algorithms can utilize detailed knowledge about building operations, environmental parameters, and user preferences to recommend optimal strategies [12, 99]. For example, a knowledge-based algorithm can be programmed to maintain acceptable indoor pollutant concentration levels. When sensors detect pollutant levels exceeding predefined thresholds, the system recommends increased ventilation to maintain air quality [100]. Moreover, knowledge-based approaches are often used to enhance other algorithms. For instance, the HeartSteps algorithm, developed to promote physical activity, integrates RL with prior knowledge derived from historical data to overcome the cold start problem typical of traditional recommender systems [14].
@@
Clustering algorithms enhance targeted and efficient recommendations by grouping similar items or users based on their attributes or behavioral characteristics using clustering techniques [101]. Unlike collaborative filtering, which relies on user ratings or behavioral similarity [102], clustering algorithms focus on identifying the inherent grouping structure within the data [103, 104], thereby significantly mitigating the cold start problem commonly encountered in collaborative filtering due to data sparsity [105]. For instance, clustering methods like K-means [106] and DBSCAN [107] can perform extensive computations during the preprocessing stage, thereby reducing the need for real-time calculations and improving the operational efficiency of recommender systems. Despite their advantages in handling large-scale datasets, clustering algorithms exhibit limitations in dynamic environments, such as building control or indoor environment optimization [108]. Clustering algorithms typically require major computations, such as determining cluster centers and assigning data points to clusters, to be completed during the preprocessing phase [109]. This means that whenever there are changes in the dataset, such as the addition of building operational data, the clustering model may need to be rerun to reflect these changes. This process is not automatic, contrasting with the real-time responsiveness of collaborative filtering [102].
@@
Deep learning-based algorithms utilize complex neural network architectures to model user preferences and item characteristics within a recommendation framework [118]. These systems employ multi-layer nonlinear processing units to learn deep representations of features from complex user behaviors, item properties, and textual descriptions, enabling them to discern intricate patterns in large-scale data that traditional algorithms might overlook [119, 120]. For instance, in [121], deep learning models were used to address significant cold start issues in collaborative filtering systems by embedding information about users and their preferences directly into the model. Moreover, deep learning techniques are frequently applied to process unstructured data, such as text, images, and audio [122]. Such as Convolutional Neural Networks (CNNs) [123, 124] are used for image recommendations, while Recurrent Neural Networks (RNNs) [125] handle sequential interaction data. Emerging technologies like attention mechanisms and Transformer models [126] have been widely adopted in deep learning-based recommender systems [119]. These models focus more on the most pertinent parts of the input data, dynamically adjusting the importance weights of different segments to optimize information processing and feature extraction [127]. For instance, in [128], the incorporation of time-aware self-attention mechanisms enhances the precision of capturing user preferences, significantly improving the performance of recommendation models.
4.2 Well-established applications of recommender systems in human-building interaction